an experienced bricklayer can construct a small wall in 3 hours. If him and his apprentice can complete the job together in 2 hours, how long does it take his apprentice to build the same wall by himself
Answer by checkley77(12844)

1/3+1/X=1/2
1/X=1/2-1/3
1/X=3/6-2/6
1/X=1/6
X=6 HOURS FOR THE APPRENTCE WORKING ALONE.
PROOF:
1/3+1/6=1/2
2/6+1/6=1/2
3/6=1/2
1/2=1/2